OpenStudy (anonymous):

Solve for x 3sinx - 2cosx^2x = 0

OpenStudy (anonymous):

3sinx -2cos²x + 2sin²x = 0 3sinx -2+4sin ²x=0 then factorise them

OpenStudy (callisto):

@sparkly16 Sorry to ask, but is this your question? \[3sinx - 2cos^2x = 0\]

OpenStudy (anonymous):

3sinx−2cos² 2x=0

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

3sinx−2cos 2x=0 tis one

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Once again... Is it (i) or (ii)? \[(i)3sinx - 2cos^2x = 0\]\[(ii)3sinx - 2cos2x = 0\]

OpenStudy (anonymous):

its (i)

OpenStudy (anonymous):

oh so it is 3six x -2 +2 sin²x then sin x =0.5?

OpenStudy (callisto):

Use the identity \[sin^2x+cos^2x=1\] \[3sinx - 2cos^2x = 0\]\[3sinx - 2(1-sin^2x) = 0\]\[3sinx - 2 + 2sin^2x = 0\]\[2sin^2x + 3sinx - 2 = 0\]\[(2sinx-1)(sinx +2)=0\]\[2sin^2x + 3sinx - 2 = 0\]\[(2sinx-1)=0 \ or \ (sinx +2)=0\]\[sinx=\frac{1}{2} \ or \ sinx = -2 (rejected)\]So, sinx = 0.5 Take arc sin to find the value of x, can you?

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (callisto):

Welcome :)